# Control of kinetic opinion dynamics in popularity-adaptive social networks

### v1

@@ 3 The controlled opinion evolution
$\partial_{t}f(v,c,t)=$ $-\alpha\partial_{v}\left(\left(\mathcal{P}[f](v,c,t)+\frac{R_{v}(c)H_{v}(\rho(v))\,\left(\,v-\tilde{v}\,\right)}{\gamma_{v}}\right)f(v,c,t\right)$ (14)
$-\beta\partial_{c}\left(\left(\frac{\mu}{2}\ln\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\lambda\,\frac{\beta}{\gamma_{c}}\,R_{c}(c)\,H_{c}(v)\right)cf(v,c,t)\right)$
$+\frac{\sigma^{2}}{2}\partial_{vv}^{2}\left(D(v)^{2}f(v,c,t)\right)+\frac{\nu^{2}}{2}\partial_{cc}^{2}\left(c^{2}f(v,c,t)\right).$
@@ 4 Numerical experiments
In this Section we present a set of numerical experiments aimed at illustrating the behavior of the proposed model and analyzing the joint evolution of contacts and opinions when the control mechanisms introduced in the previous sections are active.
For the numerical approximation of the Boltzmann-type equation (6) in the Fokker–Planck regime (10)–(14), we adopt a Monte Carlo–type strategy. In particular, we rely on an asymptotic stochastic particle method suitable for kinetic equations in the quasi-invariant limit, introducing the discrete update
$f^{n+1}=\left(1-\frac{\Delta t}{\varepsilon}\right)f^{n}+\frac{\Delta t}{\varepsilon}\,\mathcal{G}_{\varepsilon}(f^{n},f^{n}),$ (15)
where the gain operator $\mathcal{G}_{\varepsilon}$ accounts for the inflow of particles at position $(v,c)$ at time $t$ after the interaction rules (2) and (11) are applied.
The particle scheme used to simulate (15) is detailed in Algorithm 1, where we take $\varepsilon=\Delta t$ for simplicity. Further details on this class of asymptotic particle methods can be found, e.g., in [PT13].
Algorithm 1 Asymptotic particle-based scheme (Nanbu-like algorithm)
@@
Sample $\{v_{i}^{0},c_{i}^{0}\}_{i=1}^{N_{s}}$ from the initial distribution $f_{0}(v,c)$.
@@
   Compute the contact-control contributions
$\kappa_{j}^{n}=\lambda\frac{\beta}{\gamma_{c}}R_{c}(c^{n}_{j})H_{c}(v^{n}_{j}),\qquad j\in\{i,i_{*}\}.$
   Compute the opinion controls
$u_{i}^{n}$ $=-\frac{R_{v}(c_{i}^{n})H_{v}(v_{i}^{n})\left(v_{i}^{n}+\varepsilon\alpha(P(v_{i}^{n},v_{i_{*}}^{n},c_{i}^{n},c_{i_{*}}^{n})(v_{i_{*}}^{n}-v_{i}^{n}))-\tilde{v}\right)}{\gamma_{v}+\varepsilon\alpha R_{v}(c_{i}^{n})H_{v}(v_{i}^{n})},$
$u_{i_{*}}^{n}$ $=-\frac{R_{v}(c_{i_{*}}^{n})H_{v}(v_{i_{*}}^{n})\left(v_{i_{*}}^{n}+\varepsilon\alpha(P(v_{i_{*}}^{n},v_{i}^{n},c_{i_{*}}^{n},c_{i}^{n})(v_{i}^{n}-v_{i_{*}}^{n}))-\tilde{v}\right)}{\gamma_{v}+\varepsilon\alpha R_{v}(c_{i_{*}}^{n})H_{v}(v_{i_{*}}^{n})}.$
   Update contacts
$c_{j}^{n+1}=c_{j}^{n}-\varepsilon\beta\!\left(\Psi^{\varepsilon}\!\left(\frac{c_{j}^{n}}{\overline{c}}\right)+\Phi(v_{j}^{n})-\kappa_{j}^{n}\right)c_{j}^{n}+\eta_{\varepsilon,j}^{n}\,c_{j}^{n},\qquad j\in\{i,i_{*}\}.$
   Update opinions
$v_{i}^{n+1}$ $=v_{i}^{n}+\varepsilon\alpha\!\left(P(v_{i}^{n},v_{i_{*}}^{n},c_{i}^{n},c_{i_{*}}^{n})+u_{i}^{n}\right)+\sqrt{\varepsilon}\,\sigma\,D(v_{i}^{n})\,\xi_{i}^{n},$
$v_{i_{*}}^{n+1}$ $=v_{i_{*}}^{n}+\varepsilon\alpha\!\left(P(v_{i_{*}}^{n},v_{i}^{n},c_{i_{*}}^{n},c_{i}^{n})+u_{i_{*}}^{n}\right)+\sqrt{\varepsilon}\,\sigma\,D(v_{i_{*}}^{n})\,\xi_{i_{*}}^{n}.$
@@
Table 1: Summary of model parameters.
@@
Endogenous modulation $\Psi(\cdot)$ $\mu=0$, $\bar{c}=200$
@@
Cost penalization $\gamma_{c}=1$
Maximal control intensity $\lambda=1$
Activation thresholds $\alpha_{R}=\alpha_{H}=0.1$, $r=0.7$, $\rho^{\star}=0.5$
Contact activation level $c_{\min}=100$
Opinion dynamics
Binary interaction $\alpha=1$, $\delta=0.8$, $p=3$, $\tilde{v}=0.5$
Opinions noise $\xi$, $\sigma=0.1$
@@
Cost penalization $\gamma_{v}=10$
Thresholds $R_{v}(c)=H_{v}(v)=1$
In this test we investigate the role of contact and opinion control mechanisms in a structured “leader–follower” system, where leaders denote agents endowed with control authority (or externally driven dynamics), while followers are agents that evolve according to interaction rules and respond to the influence of the leaders.
We consider a population of agents, $25\%$ of which are designated as “leaders”, whereas the other represent the “mass”. The simulation runs up to a final time $T=50$ with uniform time step $\Delta t=10^{-3}$, and the all the others parameters used in this test are listed in Tab. 1.
@@ 4.1 Test 1: comparative analysis of control strategies in coupled contact–opinion dynamics
We sample $N_{s}=10^{6}$ agents to simulate the model. At the initial time $t=0$, the joint density $f_{0}(c,v)=f(c,v,0)$ is represented in Fig. 2, and is provided as follows: the leaders’ states $(c,v)$ are sampled from a continuous uniform distribution $(c,v)\sim\mathrm{Unif}\big([150,200]\times[0.4,0.6]\big),$ representing a cohesive and highly connected group with an opinion significantly different from that of the mass, which instead is initialized according to $(c,v)\sim\mathrm{Unif}\big([10,90]\times[-0.9,-0.1]\big),$ corresponding to a low-connectivity population spread over a broad range of opinions.
We examine four scenarios: (a) fully uncontrolled dynamics; (b) contact control $\kappa^{\star}$ active only for leaders; (c) opinion control $u^{\star}$ active only for leaders; (d) both controls simultaneously active on leaders.
@@
Figure 3: Test 1. Time evolution of the joint density $f(c,v,t)$ in the $(v,c)$ plane for different control strategies. Each panel shows snapshots at times $t=1,5,50$. (a) uncontrolled dynamics; (b) control on contacts; (c) control on opinions; (d) both controls active.
@@
Simulation results are reported in Fig. 3, which displays snapshots of the joint density $f(c,v,t)$ at selected times for each scenario, and in Fig. 4, illustrating the temporal evolution of the marginal distributions of contacts $h(c)$ and opinions $g(v)$.
In the uncontrolled case (a), the leader group, initially located around an extreme opinion, progressively loses contacts, thereby reducing its influence, and is eventually absorbed into the followers’ cluster, converging to a compromise opinion. In scenario (b), leaders initially lose popularity, but the contact control activates near the threshold $c_{\min}$, preventing collapse and enabling them to regain influence; as a result, the overall opinion shifts toward a compromise closer to the leaders’ initial stance. Scenario (c), with only opinion control active, shows that the loss of contacts dominates the dynamics: leaders rapidly lose influence and merge into the followers’ distribution, again converging to a compromise. Finally, in scenario (d), the combined action of both controls preserves leaders’ popularity and enables them to steer the population toward their target opinion.
In Figure 5 we report the time evolution of the mean number of contacts $m_{c}(t)$ and of the mean opinion $m_{v}(t)$ for the four scenarios (a)–(d). The plots show that when the contact control is active, namely in scenarios (b) and (d), the global mean of contacts exhibits a sustained growth over time, counteracting the natural decay produced by the uncontrolled dynamics. Moreover, in scenario (d), where both opinion and contact controls are simultaneously activated, the combined action of the two control layers drives the population’s mean opinion towards the target value promoted by the leader group, highlighting the synergistic effect of the coupled control mechanisms.
@@
In this test we study the effect of opinion control in a social system consisting of two competing leader groups, each promoting an opposite opinion, and a third uncontrolled group representing the mass of followers. The aim is to analyse how symmetric or asymmetric penalization of the opinion-control effort influences the outcome of the competition for consensus.
We consider a population of agents divided as follows: $25\%$ constitute Leader Group A, another $25\%$ constitute Leader Group B, and the remaining $50\%$ form the follower population: the “mass”. The simulation horizon is $T=50$ with time step $\Delta t=10^{-3}$.
The evolution of contacts and opinions follows the same parameters used in Test 1 as in Tab 1. The only difference concerns the penalization parameters for the opinion control, denoted here by $\gamma_{v}^{A}$ and $\gamma_{v}^{B}$ for the two leader groups, and for the target opinions $\tilde{v}_{A},\tilde{v}_{B}$.
@@
Figure 7: Test 2. Time evolution of the joint density $f(c,v,t)$ in the $(v,c)$ plane. Each panel shows snapshots at times $t=1,5,10,15,20$. (a) uncontrolled dynamics; (b) symmetric opinion control; (c) non-symmetric opinion control.
@@
Simulation results are presented in Fig. 7, showing snapshots of the joint density $f(c,v,t)$ at selected times, and in Fig. 8, illustrating the evolution of the marginal distributions of contacts $h(c)$ and opinions $g(v)$.

### v2

@@ 3 The controlled opinion evolution
$\partial_{t}{\color[rgb]{0,0,0}{f_{\ell}(v,c,t)}}=$ $-\alpha\partial_{v}\left(\left({\color[rgb]{0,0,0}{\sum_{k}\mathcal{U}_{\ell}[f_{k}](v,c,t)}}\right){\color[rgb]{0,0,0}{f_{\ell}(v,c,t)}}\right)$ (18)
$-\partial_{c}\left(\left(-\frac{\mu}{2}\ln\left(\frac{c}{\overline{c}}\right)-{\color[rgb]{0,0,0}{\beta\left(\Phi(v)-\kappa_{\ell}\right)}}\right)c{\color[rgb]{0,0,0}{f_{\ell}(v,c,t)}}\right)$
$+\frac{\sigma^{2}}{2}\partial_{vv}^{2}\left(D(v)^{2}{\color[rgb]{0,0,0}{f_{\ell}(v,c,t)}}\right)+\frac{\nu^{2}}{2}\partial_{cc}^{2}\left(c^{2}{\color[rgb]{0,0,0}{f_{\ell}(v,c,t)}}\right).$
Remark 3.1 (Structure of the optimal control term).
In the general case, $\bar{u}_{\ell}$ may depend on both $(v,c)$ and the states of the interacting agents $(v_{*},c_{*})$. For the specific choice of the control (13) we have
$\bar{u}_{\ell}=-\,\frac{R^{\ell}_{v}(c)H^{\ell}_{v}(v)\,(\,v-\tilde{v}_{\ell}\,)}{\gamma^{\ell}_{v}},$
which depends only on $(v,c)$, and (17) simplifies accordingly, reducing to
$\mathcal{U}_{\ell}[f_{k}](v,c,t)=\iint_{[-1,1]\times\mathbb{R}_{+}}P(v,v_{*},c,c_{*})(v_{*}-v)f_{k}(v_{*},c_{*},t)\,dv_{*}dc_{*}+\bar{u}_{\ell}(v,c)\rho_{k}.$
Remark 3.2 (Macroscopic moment dynamics).
The evolution of macroscopic quantities is obtained from the mesoscopic model, such as (18), by taking velocity moments of the distribution function and integrating with respect to the microscopic variables. In this case, the presence of the nonlocal operator generally prevents the closure of the resulting moment system unless an equilibrium closure is assumed. Nevertheless, by imposing suitable simplifying assumptions on the nonlinear functions, one can derive a system of equations governing the first-order moments with respect to $v$. In particular, assume that $P(v,v_{*},c,c_{*})=1$ and that, in the control term (13), the functions $R_{v}(c)$ and $H_{v}(v)$ are both identically equal to $1$. Then, multiplying (18) by $v$ and integrating with respect to $v$, we obtain the first moment equation as follows
$\frac{d}{dt}m_{\ell,v}(t)$ $=\alpha\left(m_{v}(t)-m_{\ell,v}(t)\right)+\frac{\alpha}{\gamma_{v}^{\ell}}\left(\tilde{v}_{\ell}-m_{\ell,v}(t)\right)$ (19)
resulting in a system where $m_{v}(t):=\sum_{k}\rho_{k}m_{k,v}(t)$ denotes the average opinion of the whole population. Furthermore, we observe that by assuming the same penalization parameter $\gamma_{v}^{\ell}\equiv\gamma_{v}$, multiplying by $\rho_{\ell}$, and summing over $\ell$ in (20), we can write an equation for $m_{v}(t)$ as follows
$\frac{d}{dt}m_{v}(t)$ $=\frac{\alpha}{\gamma_{v}}\left(\sum_{\ell}\rho_{\ell}\tilde{v}_{\ell}-m_{v}(t)\right),$ (20)
which corresponds to a relaxation toward the weighted average of the desired target opinions $\tilde{v}_{\ell}$ of the different populations.
@@ 4 Numerical experiments
In this Section we present a set of numerical experiments aimed at illustrating the behavior of the proposed model and analyzing the joint evolution of contacts and opinions in the presence of the control mechanisms introduced in the previous sections.
In the multi-population setting described above, the state of the system is given by a family of distribution functions $\{f_{\ell}(v,c,t)\}_{\ell}$, one for each population, whose evolution is governed by a system of Boltzmann-type equations (7) coupled through cross-interaction operators. In the Fokker–Planck regime (18), we approximate this system by means of a Monte Carlo–type strategy based on an asymptotic stochastic particle method, suitable for kinetic equations in the quasi-invariant limit.
More precisely, denoting by $f_{\ell}^{n}$ the numerical approximation of $f_{\ell}$ at time $t^{n}$, we consider the discrete update
$f_{\ell}^{n+1}=\left(1-\frac{\Delta t}{\varepsilon}\right)f_{\ell}^{n}+\frac{\Delta t}{\varepsilon}\,\sum_{k}\mathcal{G}_{\ell,\varepsilon}\!\left(f_{\ell}^{n},f_{k}^{n}\right),$ (21)
where the gain operators $\mathcal{G}_{\ell,\varepsilon}$ account for the inflow of particles in population $\ell$ at $(v,c)$ due to interactions with agents of population $k$, according to the microscopic rules (3) and (12). The corresponding particle scheme is detailed in Algorithm 1, where, for simplicity, we take $\varepsilon=\Delta t$. Further details on this class of asymptotic particle methods for kinetic equations in the quasi-invariant limit can be found, e.g., in [34].
Algorithm 1 Asymptotic particle-based scheme (Nanbu-like algorithm) with multiple populations
@@
Sample $\{v_{i}^{0},c_{i}^{0},\ell_{i}\}_{i=1}^{N_{s}}$ from the initial multi-population distribution $\{f^{0}_{\ell}(v,c)\}$.
@@
   Compute the contact controls $\kappa_{i}^{n}=(\kappa_{\ell})_{i}^{n}$ and $\kappa_{i_{*}}^{n}=(\kappa_{\ell_{*}})_{i_{*}}^{n}$.
$\kappa_{i}^{n}=\frac{\lambda^{\ell}}{\gamma^{\ell}_{c}}R^{\ell}_{c}(c^{n}_{i})H^{\ell}_{c}(v^{n}_{i}),\qquad\kappa_{i_{*}}^{n}=\frac{\lambda^{\ell_{*}}}{\gamma^{\ell_{*}}_{c}}R^{\ell_{*}}_{c}(c^{n}_{i_{*}})H^{\ell_{*}}_{c}(v^{n}_{i_{*}}).$
   Compute the opinion controls $u_{i}^{n}=(u_{\ell})_{i}^{n}$ and $u_{i_{*}}^{n}=(u_{\ell_{*}})_{i_{*}}^{n}$.
$u_{i}^{n}$ $=-\frac{R_{v}^{\ell}(c_{i}^{n})H_{v}^{\ell}(v_{i}^{n})\left(v_{i}^{n}+\varepsilon\alpha P(v_{i}^{n},v_{i_{*}}^{n},c_{i}^{n},c_{i_{*}}^{n})(v_{i_{*}}^{n}-v_{i}^{n})-\tilde{v}^{\ell}\right)}{\gamma_{v}^{\ell}+\varepsilon\alpha R_{v}^{\ell}(c_{i}^{n})H_{v}^{\ell}(v_{i}^{n})},$
$u_{i_{*}}^{n}$ $=-\frac{R_{v}^{\ell_{*}}(c_{i_{*}}^{n})H_{v}^{\ell_{*}}(v_{i_{*}}^{n})\left(v_{i_{*}}^{n}+\varepsilon\alpha P(v_{i_{*}}^{n},v_{i}^{n},c_{i_{*}}^{n},c_{i}^{n})(v_{i}^{n}-v_{i_{*}}^{n})-\tilde{v}^{\ell_{*}}\right)}{\gamma_{v}^{\ell_{*}}+\varepsilon\alpha R_{v}^{\ell_{*}}(c_{i_{*}}^{n})H_{v}^{\ell_{*}}(v_{i_{*}}^{n})}.$
   Update contacts.
$c_{j}^{n+1}=c_{j}^{n}-\Psi^{\varepsilon}\!\left(\frac{c_{j}^{n}}{\overline{c}}\right)-\varepsilon\beta\!\left(\Phi(v_{j}^{n})-\kappa_{j}^{n}\right)c_{j}^{n}+\eta_{\varepsilon,j}^{n}\,c_{j}^{n},\qquad j\in\{i,i_{*}\}.$
   Update opinions.
$v_{i}^{n+1}$ $=v_{i}^{n}+\varepsilon\alpha\!\left(P(v_{i}^{n},v_{i_{*}}^{n},c_{i}^{n},c_{i_{*}}^{n})(v_{i_{*}}^{n}-v_{i}^{n})+u_{i}^{n}\right)+\sqrt{\varepsilon}\,\sigma\,D(v_{i}^{n})\,\xi_{i}^{n},$
$v_{i_{*}}^{n+1}$ $=v_{i_{*}}^{n}+\varepsilon\alpha\!\left(P(v_{i_{*}}^{n},v_{i}^{n},c_{i_{*}}^{n},c_{i}^{n})(v_{i}^{n}-v_{i_{*}}^{n})+u_{i_{*}}^{n}\right)+\sqrt{\varepsilon}\,\sigma\,D(v_{i_{*}}^{n})\,\xi_{i_{*}}^{n}.$
@@ 4.1 Test 1: comparative analysis of control strategies in coupled contact–opinion dynamics
In this test we investigate the role of contact and opinion control mechanisms for two populations of agents, structured as a “leader–follower” system, where leaders denote agents endowed with control authority (or externally driven dynamics), while followers are agents that evolve according to interaction rules and respond to the influence of the leaders. In particular, $25\%$ of the agents are designated as “leaders” (population $\ell=L$), whereas the others represent the “followers” (population $\ell=F$).
Table 1: Summary of model parameters, the superscripts in the control terms correspond to the population $\ell=L$ for leaders and $\ell=F$ for followers.
@@
Endogenous modulation $\Psi(\cdot)$ $\mu=0.25$, $\bar{c}=100$
@@
Opinion dynamics
Binary interaction $\alpha=1$, $\delta=0.8$, $p=3$
Opinions noise $\xi$, $\sigma=0.1$
@@
Cost penalization $\gamma^{L}_{c}=1$
Maximal control intensity $\lambda^{F}=0$, $\lambda^{L}=1$ (when activated)
Activation thresholds $\alpha^{L}_{R}=\alpha^{L}_{H}=0.1$, $r^{L}=0.7$, $\rho_{\star}^{L}=0.5$
Contact activation level $c^{L}_{\min}=150$
@@
Thresholds $R^{F}_{v}(c)=H^{F}_{v}(v)=0$, $R^{L}_{v}(c)=H^{L}_{v}(v)=1$
Target opinion $\tilde{v}^{L}=0.5$
Cost penalization $\gamma_{v}^{L}=10$
@@
We sample $N_{s}=10^{6}$ agents to simulate the model. At the initial time $t=0$, the total joint density $f_{0}(c,v)=f(c,v,0)$ is represented in Fig. 2, and is provided as follows: the leaders’ states $(c,v)$ are sampled from a continuous uniform distribution $(c,v)\sim\mathrm{Unif}\big([200,250]\times[0.4,0.6]\big),$ representing a cohesive and highly connected group with an opinion significantly different from that of the mass, which instead is initialized according to $(c,v)\sim\mathrm{Unif}\big([10,90]\times[-0.9,-0.1]\big),$ corresponding to a low-connectivity population spread over a broad range of opinions.
The simulation runs up to a final time $T=50$ with uniform time step $\Delta t=10^{-3}$.
In this test the followers are always uncontrolled, while for the leaders we switch on/off the contact control $\kappa^{\star}$ and the opinion control $u^{\star}$, obtaining the four scenarios (a)–(d) described below: (a) fully uncontrolled dynamics for both populations; (b) contact control $\kappa^{\star}$ active only on the leader population; (c) opinion control $u^{\star}$ active only on the leader population; (d) both controls simultaneously active on the leader population. In all cases, the follower population evolves without control.
All numerical parameters are listed in Table 1; only the activation of the control terms changes across scenarios.
@@
[…]
@@
[…]
In the uncontrolled case (a), the leader group, initially located around an extreme opinion, progressively loses contacts, thereby reducing its influence, and is eventually absorbed into the followers’ cluster, converging to a compromise opinion. In scenario (b), leaders initially lose popularity, but the contact control activates near the threshold $c^{L}_{\min}$, preventing collapse and enabling them to regain influence; as a result, the overall opinion shifts toward a compromise closer to the leaders’ initial stance. Scenario (c), with only opinion control active, shows that the loss of contacts dominates the dynamics: leaders rapidly lose influence and merge into the followers’ distribution, again converging to a compromise. Finally, in scenario (d), the combined action of both controls preserves leaders’ popularity and enables them to steer the population toward their target opinion.
In Figure 5 we report the time evolution of the mean number of contacts $m_{c}(t)$ and of the mean opinion $m_{v}(t)$ for the four scenarios (a)–(d). The plots show that when the contact control is active, namely in scenarios (b) and (d), a larger and more persistent increase in the global mean of contacts is displayed, compared to the remaining cases. Moreover, in scenario (d), where both opinion and contact controls are simultaneously activated, the combined action of the two control layers drives the population’s mean opinion towards the target value promoted by the leader group, highlighting the synergistic effect of the coupled control mechanisms.
To address the effect of the fraction of leaders, we repeated scenario (d) varying the initial proportion of leaders from 1% up to 25%, in a larger temporal interval with $T=100$. We observe that as the fraction of leaders increases, the attraction of the population towards the target opinion promoted by the leaders becomes stronger. Interestingly, at the initial stages, the opinions of the leaders tend to slightly diverge from their own mean target opinion, while the followers’ opinions progressively align with those of the leaders, the rate of alignment depending on the strength of the attraction mechanism. These dynamics are illustrated in Fig. 6.
@@
Figure 6: Test 1. Time evolution of the mean opinion $m_{v}(t)$ over the interval $[0,100]$ for Leaders and Followers in scenario (d). Each color corresponds to a different fraction of leaders (1%, 5%, 15%, 25%). Line styles distinguish the two populations: continuous lines represent Leaders, while dotted lines represent Followers. The horizontal dashed line indicates the target opinion at $0.5$.
@@
In this test we study the effect of opinion control in a social system consisting of two competing leader groups, each promoting an opposite opinion, and a third uncontrolled group representing the mass of followers. In particular, we analyze how symmetric or asymmetric penalization of the opinion-control effort influences the outcome of the competition for consensus.
We consider three populations of agents divided as follows: $25\%$ constitute Leader Group A ($\ell=A$), another $25\%$ constitute Leader Group B ($\ell=B$), and the remaining $50\%$ form the “mass” ($\ell=M$). The simulation horizon is $T=50$ with time step $\Delta t=10^{-3}$.
The evolution of contacts and opinions follows the same parameters used in Test 1 as in Tab 1. The only difference concerns the penalization parameters for the opinion control, indicated here by $\gamma_{v}^{A}$ and $\gamma_{v}^{B}$ for the two leader populations and for the target opinions $\tilde{v}_{A},\tilde{v}_{B}$.
@@
[…]
@@
[…]